# Temporal Wasserstein Imputation: Versatile Missing Data Imputation for Time Series

### v2

@@ 5.1 Linear and nonlinear univariate time series
In this subsection, we consider the following data-generating processes (DGPs) for univariate time series, which can exhibit linear or nonlinear dynamic patterns. In the sequel, $\{\epsilon_{t}\}$ and $\{\epsilon_{t,j}\}$ for $j=1,2,\ldots$ denote i.i.d. sequences of standard Gaussian variables.
@@
$x_{t}=\left\{\begin{array}[]{cc}\phi_{1}x_{t-1}+\epsilon_{t},&x_{t-1}\leq\tau\\ \phi_{2}x_{t-1}+0.5\epsilon_{t},&x_{t-1}>\tau\end{array}\right.,$
where $(\phi_{1},\phi_{2},\tau)=(-2,0.7,1)$
@@
DGPs 1 and 2 are well-known linear time series models. Note that DGP 2 admits an AR($\infty$) representation, so any finite-dimensional marginal distribution is inadequate to fully characterize the distribution of the underlying process. DGP 3 is a celebrated threshold AR model which has found many applications in modeling nonlinear dynamics in the econometrics and statistics literature (Montgomery et al.,, 1998; Li and Ling,, 2012; Tsay and Chen,, 2018). DGP 4 generates an I(1) series which typically produces a highly persistent stochastic trend. Finally, DGP 5 features a deterministic cyclic trend with non-standard periods. Specifically, it includes two sinusoidal components with periods of 8.70 and 11.77, respectively.
@@
Three measures of performance are considered. First, we report the Wasserstein distance (of order 2) between the empirical 3-dimensional marginal distributions of the imputed series and the full data. Specifically, let $\{\hat{w}_{t}^{(i)}\}$ and $\{x_{t}^{(i)}\}$ be the imputed series and the original series in simulation $i$, respectively, for $i=1,2,\ldots,1000$. We report
$\frac{1}{1000}\sum_{i=1}^{1000}\mathcal{W}_{2}\left(\frac{1}{n-3}\sum_{t=2}^{n-1}\delta_{(\hat{w}_{t}^{(i)},\hat{w}_{t-1}^{(i)},\hat{w}_{t-2}^{(i)})},\frac{1}{n-3}\sum_{t=2}^{n-1}\delta_{({x}_{t}^{(i)},{x}_{t-1}^{(i)},{x}_{t-2}^{(i)})}\right),$ (14)
@@
as a measure of how well the imputed series approximates the original series. Second, for DGP 14, the model parameters are estimated using the imputed series, and their root mean square errors (RMSEs) are reported. Third, we compare the RMSEs of the estimated autocovariance functions (ACFs). However, due to space constraints, we relegate the comparison of ACFs to Section S3.1 in the supplementary material.
@@
The Wasserstein distance between the imputed and the original marginal distributions is reported in Table 1. Across all five DGPs employed, the smallest loss is consistently achieved by TWI-type methods, indicating that TWI effectively approximates the distributional properties of the underlying series. While other methods may yield good imputations for some specific DGPs, the result is not consistent across all processes. For instance, linear and spline interpolations show low Wasserstein loss for DGP 1, but their performance is less competitive when applied to other models. In contrast, TWI methods, carefully initialized, consistently deliver solid performance across models and missing patterns. For the nonstationary Model 4 (I(1)), we first obtain an imputation $\{\hat{w}_{t}\}$ and consider the differenced series $\nabla\hat{w}_{t}=\hat{w}_{t}-\hat{w}_{t-1}$. Table 1 shows the Wasserstein distance between the marginal distributions implied by $\{\nabla\hat{w}_{t}\}$ and $\{\nabla x_{t}\}$. In this case, TWI consistently yields the lowest Wasserstein loss and greatly improves from its initialization.
@@
Table 1 also shows TWI typically improves the Wasserstein loss from its initialization, often significantly. For example, for DGP 3 (TAR), TWI reduces the Wasserstein loss from Kalman smoothing initialization by approximately 45% under missing pattern I (from 1.38 to 0.74). Similarly, for DGP 5 (CYC), TWI reduces the Wasserstein loss by more than 50% from linear initialization (from 1.96 to 0.79). Figure 2 plots $\hat{w}_{t}$ against $\hat{w}_{t-1}$ when either or both are imputed under DGP 3, alongside $(x_{t},x_{t-1})$ of the original data. While Kalman smoothing introduces substantial distortion in the marginal distribution, TWI, initialized by Kalman smoothing, significantly corrects this bias, producing imputations that capture the nonlinearity in the ground truth. Note that Kalman smoothing is done via an ARIMA state-space representation; hence the model is misspecified if the underlying process is nonlinear. Nevertheless, it already serves as a good initialization of TWI.
@@
Table 1: Wasserstein distance between the empirical marginal distributions of the imputed series and those of the full data, averaged over 1000 simulations. TWIlin denotes temporal Wasserstein imputation using linear interpolation as initialization and TWIKal denotes temporal Wasserstein imputation with Kalman smoothing as initialization. Similar notations are used for $k$-TWI. For DGP 4 (I(1)), the loss is computed using the first-differenced series. The smallest value for each model and missing pattern is marked in boldface.
Model Linear Spline Kalman ScalarF TWIlin $k$-TWIlin TWIKal $k$-TWIKal
@@
AR 0.41 0.41 0.42 0.50 0.40 0.44 0.41 0.44
ARMA 0.48 0.58 0.47 0.48 0.40 0.38 0.40 0.38
TAR 1.12 1.51 1.38 1.13 0.96 0.81 0.74 0.63
I(1) 0.71 0.75 0.62 4.38 0.58 0.53 0.54 0.50
CYC 1.96 0.82 0.77 1.97 0.79 0.77 0.60 0.70
NLVAR 2.98 3.42 3.01 2.94 2.25 2.14 2.19 2.12
AL (×$\times 10$) 0.70 1.00 4.17 0.55 0.43 0.37 0.38 0.35
@@
AR 0.44 0.59 0.50 0.76 0.39 0.44 0.43 0.45
ARMA 0.47 1.29 0.51 0.53 0.36 0.34 0.40 0.35
TAR 1.04 3.11 1.25 0.99 0.84 0.73 0.76 0.61
I(1) 0.67 0.65 0.60 8.35 0.51 0.45 0.49 0.44
CYC 2.58 1.81 0.77 3.79 2.62 1.60 0.62 0.72
NLVAR 3.00 6.07 3.33 3.67 2.13 2.02 2.31 2.04
AL (×$\times 10$) 0.56 1.83 3.81 0.46 0.36 0.33 0.40 0.34
@@
Figure 2: Scatter plot of $(x_{t-1},x_{t})$ of the original data and $(\hat{w}_{t-1},\hat{w}_{t})$ of the imputed series, when the data are generated from Model 3 (TAR).
Next, we turn to the model parameters estimated from the imputed series. Table 2 presents the RMSEs for each model parameter. For DGP 1 (AR) and DGP 2 (ARMA), Kalman smoothing and the scalar filter yield satisfactory estimates among the benchmarks, with TWI offering only limited improvements. For the nonlinear DGP 3 (TAR), the results are drastically different. TWI produces the most accurate parameter estimates and can achieve more than 50% reduction in estimation errors compared to linear interpolation and Kalman smoothing. These estimation results demonstrate TWI’s strength in learning the underlying nonlinear dynamics. For DGP 4 (I(1)), TWI either outperforms or stays on par with the best benchmark, with Kalman filter being a strong initialization.
Table 2: Estimation root mean square errors for parameter estimation errors. The results are averaged over 1000 simulations. See Table 1 for notations.
Model Linear Spline Kalman ScalarF TWIlin $k$-TWIlin TWIKal $k$-TWIKal
@@ 5.2 Multivariate time series
In this subsection, we consider multivariate nonlinear time series models.
@@
$x_{t,1}=\phi_{11}x_{t-1,1}+\phi_{12}\mathfrak{s}(3x_{t-1,2})+0.25\epsilon_{t,1}\quad x_{t,2}=\phi_{21}x_{t-1,1}+\phi_{22}x_{t-1,2}+3\epsilon_{t,2}$
where $\mathfrak{s}(z)=1/(1+\exp(-z))-0.5$ is the centered sigmoid function, and $(\phi_{11},\phi_{12},\phi_{21},\phi_{22})=(0.3,8,0,0.4)$.
DGP 6 is a VAR model with a nonlinear term. In particular, the impact of $x_{t-1,2}$ on $x_{t,1}$ is nonlinear and is modeled by a sigmoid function. As depicted in Figure 3, TWI successfully captures the underlying nonlinear dynamics, whereas other imputation methods fail to preserve meaningful relationships. This is validated by the results shown in Table 1, where TWI achieves the lowest Wasserstein distance. Next, we assess the effectiveness of the imputed series in estimating the model parameters $(\phi_{11},\phi_{12},\phi_{21},\phi_{22})$, assuming the nonlinear functional form is known. The root mean square estimation errors are documented in Table 2. Compared to the benchmarks, TWI provides considerable error reduction in the parameter associated with the nonlinear effect, $\phi_{12}$. It attains more than 50% reduction in RMSE compared to the scalar filter, the best-performing benchmark.
We also consider a challenging multivariate nonlinear model, which generates the so-called compositional time series. The data are generated from the additive logistic (AL) model of Brunsdon and Smith, (1998). Due to space constraints, we present its simulation results and discussion in Section S3.2 of the supplementary material.
@@
Figure 3: Scatter plots of $(x_{t-1,i},x_{t,j})$ of the original data and $(\hat{w}_{t-1,i},\hat{w}_{t,j})$ of the imputed series for $i,j\in\{1,2\}$, when the data are generated from Model 6 under missing pattern II.
@@ 6 Groundwater data application
In this section, we apply the proposed TWI to a groundwater dataset from Taiwan, which was collected and analyzed by Hsu et al., (2020). Given the island’s steep topography and varying rainfall distribution, groundwater management is of considerable importance for Taiwan, particularly as it is a critical water source. However, the statistical analysis of groundwater levels is challenged by frequent missing data, especially from deeper wells.
@@
We aim to study the seasonal groundwater variations, for which we first transform the hourly data from 352 sites between October 1992 and August 2020 into monthly records. For each site, monthly groundwater level is calculated as the average of the observations in a given month, with months classified as missing if more than one-third of the observations are missing. Our analysis focuses on sites with a less-than-median amount of missing data, reducing the sample to 176 sites. Figure 4 shows the monthly groundwater levels for these sites, after subtracting the mean and normalizing by the standard deviation. We make two observations. First, many series exhibit strong seasonality due to the rainfall pattern in Taiwan. Monsoon and typhoons increase rainfall in summer, while winters are drier. Second, there are many outlying values and anomalies in the observed series, including abrupt changes in groundwater levels that are more than 5 standard deviations from the sample mean. We also treat such extreme values as missing data.
@@
Figure 4: Time plot of the groundwater levels of the 176 sites under study. Each series is normalized by subtracting its mean and dividing by its standard deviation.
During the period from October, 1992 to August, 2020, few months recorded valid observations at all sites. Figure 5 shows the fluctuations in the percentage of sites with missing data over time. Although missing data are mostly prevalent in the early years, there are some time during which the percentage surges, such as between 2003 and 2005. In addition, the percentage is gradually increasing in more recent years. Thus, one may lose much information if the analysis is restricted to sites with relatively complete records.
We consider two downstream tasks using the imputed series. First, following Hsu et al., (2020), we decompose the de-trended time series using the independent component analysis (ICA), for which we apply the fastICA algorithm of Hyvärinen and Oja, (2000). Because direct application of ICA becomes infeasible when the series are severely corrupted by missing data, the analysis of Hsu et al., (2020) is confined to a relatively short time span. The second downstream task considers autoregressive modeling, which is presented in Section S3.3 in the supplementary material. In implementation, we chose $n_{1}=160$ and $p=6$ for TWI, and for $k$-TWI we use $n_{1}\in\{160,225,110\}$.
Figure 5: Monthly percentage of sites recording a missing value.
Figure 6 shows the leading two independent components (ICs) estimated using the imputed series. ICs estimated from series imputed by Kalman smoothing have peaks around August-September (red shaded area) and May-June (blue shaded area), which was also observed in Hsu et al., (2020). These peaks are attributed to the dual rain seasons in Taiwan, which involve typhoons in August-September and the Meiyu period from mid-May to mid-June. ICs estimated from TWI-imputed series with Kalman initialization largely preserve these seasonal peaks, but with improved performance before 2002. Recall that missing data are prevalent during this period (Figure 5). Especially for the first IC, those computed from TWI-imputed series have clear peaks around August-September with a more consistent amplitude. On the other hand, the ICs computed from the linearly interpolated series could not effectively isolate and identify the Meiyu season, with peaks occurring slightly earlier than expected. Nonetheless, ICs estimated from TWI-imputed series with linear interpolation substantially alleviates this shortcoming and produced more interpretable results. ICs computed from spline-interpolated series also identifies the two rain seasons, albeit they show more rugged trajectories instead of clean periodic patterns. For the scalar filter, the resulting ICs similarly appear noisy, and the Meiyu season is not well-isolated, showing that the imputation induces distortions in the dynamic pattern.
Figure 6: Independent components estimated using the imputed series. The dashed red lines in each plot mark January 2002. The red and blue shaded areas correspond to August-September (typhoon season) and May-June (Meiyu season), respectively.
Overall, TWI took a more holistic view which accounts for the underlying dynamics, such as complex seasonality, and produced interpretable results. In Section S3.3 of the supplementary material, we conduct further experiments to show that the marginal distributions of TWI-imputed series, in terms of AR modeling, are also more consistent across time compared to other methods. These results show TWI performs reliably in practice and is a useful method in handling missing data in time series.

### v3

@@ 5.1 Linear and nonlinear univariate time series
We first consider the following data-generating processes (DGPs). In the sequel, $\{\epsilon_{t}\}$ and $\{\epsilon_{t,j}\}$, for $j=1,2,\ldots$, denote i.i.d. sequences of standard Gaussian variables.
@@
If $x_{t-1}\leq\tau$, then $x_{t}=\phi_{1}x_{t-1}+\epsilon_{t}$. Otherwise, $x_{t}=\phi_{2}x_{t-1}+0.5\epsilon_{t}$. The parameters $(\phi_{1},\phi_{2},\tau)=(-2,0.7,1)$.
@@
DGPs 1 and 2 are well-known stochastic time series models. Note that DGP 2 admits an AR($\infty$) representation, so any finite-dimensional marginal distribution is inadequate to fully characterize the distribution of the underlying process. DGP 3 is a celebrated threshold AR model which has found many applications in modeling nonlinear dynamics in the econometrics and statistics literature (Montgomery et al.,, 1998; Li and Ling,, 2012; Tsay and Chen,, 2018). DGP 4 generates an I(1) series which typically produces a highly persistent stochastic trend. Finally, DGP 5 features a deterministic cyclic trend with non-standard periods. Specifically, it includes two sinusoidal components with periods of 8.70 and 11.77, respectively.
@@
Three measures of performance are considered. First, we report the Wasserstein distance between the empirical 3-dimensional marginal distributions of the imputed series and the full data. Specifically, let $\{\hat{w}_{t}^{(i)}\}$ and $\{\tilde{x}_{t}^{(i)}\}$ be the imputed series and the original series in simulation $i$, respectively, for $i=1,2,\ldots,1000$. We report
$\frac{1}{1000}\sum_{i=1}^{1000}\mathcal{W}_{2}\left(\frac{1}{n-3}\sum_{t=2}^{n-1}\delta_{(\hat{w}_{t}^{(i)},\hat{w}_{t-1}^{(i)},\hat{w}_{t-2}^{(i)})},\frac{1}{n-3}\sum_{t=2}^{n-1}\delta_{(\tilde{x}_{t}^{(i)},\tilde{x}_{t-1}^{(i)},\tilde{x}_{t-2}^{(i)})}\right),$ (14)
@@
as a measure of how well the imputed series approximates the original series. Second, for DGP 14, the model parameters are estimated using the imputed series, and the estimation root mean square errors (RMSEs) are reported. Third, we compare the RMSEs of the estimated autocovariance functions (ACFs). However, due to space constraints, the results for the ACFs are moved to Section S3.1 in the supplementary material.
@@
The Wasserstein distance between the imputed and the original marginal distributions is reported in Table 1. Across all five DGPs employed, the smallest loss is consistently achieved by TWI-type methods, indicating that TWI effectively approximates the distributional properties of the underlying series. While other methods may yield good imputations for some specific DGPs, they are not consistent across all processes. For instance, linear interpolations show low Wasserstein loss for DGP 1 (AR), but it is less competitive when applied to other models. Similarly, Kalman smoothing and iSSA work well for DGP 5 (CYC), but both yield large Wasserstein loss for DGP 3 (TAR). In contrast, TWI methods, carefully initialized, consistently deliver solid performance across models and missing patterns. For the nonstationary Model 4 (I(1)), Table 1 shows the Wasserstein distance between the marginal distributions of the differenced series $\{\nabla\hat{w}_{t}\}$ and $\{\nabla x_{t}\}$. TWI again yields the lowest Wasserstein loss.
@@
Table 1 also shows TWI often significantly improves the Wasserstein loss from its initialization. For example, for DGP 3 (TAR), TWI reduces the Wasserstein loss from Kalman smoothing initialization by approximately 45% under missing pattern I (from 1.38 to 0.74). Similarly, for DGP 5 (CYC), TWI reduces the Wasserstein loss by more than 50% from linear initialization (from 1.96 to 0.79). Figure 1 plots $\hat{w}_{t}$ against $\hat{w}_{t-1}$ when either or both are imputed under DGP 3, alongside $(\tilde{x}_{t},\tilde{x}_{t-1})$ of the original data. While Kalman smoothing introduces substantial distortion in the marginal distribution, TWI, initialized by Kalman smoothing, significantly corrects this bias, producing imputations that capture the nonlinearity in the data generating process. Note that Kalman smoothing is based on an ARIMA state-space representation; hence the model is misspecified. Nevertheless, it already serves as a good initialization for TWI.
@@
Table 1: Wasserstein distance between the empirical marginal distributions of the imputed series and those of the full data, averaged over 1000 simulations. iSSA denotes gap filling by the iterative singular spectrum analysis. TWIlin denotes temporal Wasserstein imputation using linear interpolation as initialization and TWIKal denotes temporal Wasserstein imputation with Kalman smoothing as initialization. Similar notations are used for $k$-TWI. For DGP 4 (I(1)), the loss is computed using the first-differenced series. The smallest value for each model and missing pattern is marked in boldface.
Model Linear iSSA Kalman ScalarF TWIlin $k$-TWIlin TWIKal $k$-TWIKal
@@
AR 0.41 0.40 0.42 0.50 0.40 0.44 0.41 0.44
ARMA 0.48 0.51 0.47 0.48 0.40 0.38 0.40 0.38
TAR 1.12 1.29 1.38 1.13 0.96 0.81 0.74 0.63
I(1) 0.71 0.68 0.62 4.38 0.58 0.53 0.54 0.50
CYC 1.96 0.61 0.77 1.97 0.79 0.77 0.60 0.70
NLVAR 2.98 3.04 3.01 2.94 2.25 2.14 2.19 2.12
AL ($\times 10$) 0.70 0.85 4.17 0.55 0.43 0.37 0.38 0.35
@@
AR 0.44 0.43 0.50 0.76 0.39 0.44 0.43 0.45
ARMA 0.47 0.69 0.51 0.53 0.36 0.34 0.40 0.35
TAR 1.04 1.58 1.25 0.99 0.84 0.73 0.76 0.61
I(1) 0.67 0.60 0.60 8.35 0.51 0.45 0.49 0.44
CYC 2.58 1.47 0.77 3.79 2.62 1.60 0.62 0.72
NLVAR 3.00 3.17 3.33 3.67 2.13 2.02 2.31 2.04
AL ($\times 10$) 0.56 1.04 3.81 0.46 0.36 0.33 0.40 0.34
@@
Figure 1: Scatter plots of $(x_{t-1},x_{t})$ of the original data and $(\hat{w}_{t-1},\hat{w}_{t})$ of the imputed series, when the data are generated from Model 3 (TAR).
Next, we turn to the model parameters estimated from the imputed series. Table 2 presents the RMSEs for each model parameter. For DGP 1 (AR) and DGP 2 (ARMA), Kalman smoothing and the scalar filter yield satisfactory estimates among the benchmarks, with TWI offering only limited improvements. For the nonlinear DGP 3 (TAR), the results are drastically different. TWI produces the most accurate parameter estimates and can achieve more than 50% reduction in estimation errors compared to linear interpolation and Kalman smoothing. These estimation results demonstrate TWI’s strength in learning the underlying nonlinear dynamics. For DGP 4 (I(1)), TWI either outperforms or stays on par with the best benchmark, with Kalman smoothing being a strong initialization.
Table 2: Estimation root mean squared errors for parameter estimation errors. The results are averaged over 1000 simulations. See Table 1 for notations.
Model Linear iSSA Kalman ScalarF TWIlin $k$-TWIlin TWIKal $k$-TWIKal
@@ 5.2 Multivariate time series
We consider next models for multivariate nonlinear time series.
@@
Let $\mathfrak{s}(z)=1/(1+\exp(-z))-0.5$ be the centered sigmoid function. The data are generated as
$x_{t,1}=\phi_{11}x_{t-1,1}+\phi_{12}\mathfrak{s}(3x_{t-1,2})+0.25\epsilon_{t,1},\quad x_{t,2}=\phi_{21}x_{t-1,1}+\phi_{22}x_{t-1,2}+3\epsilon_{t,2},$
where $(\phi_{11},\phi_{12},\phi_{21},\phi_{22})=(0.3,8,0,0.4)$.
DGP 6 is a VAR model with a nonlinear term. In particular, the impact of $x_{t-1,2}$ on $x_{t,1}$ is nonlinear and is modeled by a sigmoid function. As depicted in Figure 2, TWI successfully captures the underlying nonlinear dynamics, whereas other imputation methods fail to preserve meaningful relationships. This is validated by the results shown in Table 1, where TWI achieves the lowest Wasserstein distance. Moreover, as shown in Table 2, TWI provides considerable error reduction in the parameter associated with the nonlinear effect, $\phi_{12}$, compared to the benchmarks. It attains more than 50% reduction in RMSE compared to the scalar filter, the best-performing benchmark.
We also consider a challenging multivariate nonlinear model, which generates the so-called compositional time series. The data are generated from the additive logistic (AL) model of Brunsdon and Smith, (1998). Due to space constraints, we present the simulation results and discussion in Section S3.2 of the supplementary material.
@@
Figure 2: Scatter plots of $(x_{t-1,i},x_{t,j})$ of the original data and $(\hat{w}_{t-1,i},\hat{w}_{t,j})$ of the imputed series for $i,j\in\{1,2\}$, when the data are generated from Model 6 under missing pattern $I$.
@@ 6 Groundwater data application
In this section, we apply the proposed TWI to a Taiwan groundwater dataset, which was collected and analyzed by Hsu et al., (2020). Because of the island’s steep topography and uneven rainfall distribution, groundwater management is of considerable importance, particularly as it is a critical water source. However, the statistical analysis of groundwater levels is plagued by frequent missing data, especially from deeper wells.
@@
To study the seasonal groundwater variations, we first transform the hourly data between October 1992 and August 2020 into monthly records. For each site, monthly groundwater level is calculated as the average in a given month, with months classified as missing if more than one-third of the observations are missing. There are many outlying values in the series, including abrupt changes in groundwater levels that are more than 5 standard deviations away from the sample mean. We also treat such extreme values as missing data. For demonstration, our analysis focuses on the sites with less than 10 missing values, reducing the sample to 15 sites. As shown in Figure 3, these series exhibit seasonality and structural changes, which are quite representative of the dataset and are common real-world features.
@@
Figure 3: Time plots of the groundwater levels of the 15 sites under study. Each series is normalized by subtracting its mean and dividing by its standard deviation.
In the following, we randomly omit $\lfloor\rho n\rfloor$ observations to create missing data, where $\rho\in\{0.1,0.2,0.3,0.4,0.5,0.6\}$ is the missing ratio, $n=308$ is the number of observations (including missing values) in the period under study, and $\lfloor x\rfloor$ is the largest integer $\leq x$ for $x>0$. However, the last 36 observations of each series are left intact because otherwise, the SSA could not be implemented due to lack of intact columns in the trajectory matrix. Since TWI performs best with stationary series, we first use kernel smoothing (estimated with missing data) to remove the trends, and then apply ($k$-)TWI to the residuals. In addition to the methods employed in Section 5, we also compare with imputing the missing values using the sample mean. Due to space constraints, the implementation details of the methods are provided in Section S3.3 of the supplementary material.
We consider two downstream tasks. First, the imputed series are used to estimate some summary statistics, including the autocovariances, autocorrelations, and partial autocorrelations. Specifically, let $\gamma_{j}(h)$ be the statistic at the $h$-th lag obtained using the $j$-th series (without the artificial missing data), and $\hat{\gamma}_{j}(h)$ be the corresponding estimates from the imputed series. Table 3 reports the mean absolute error (MAE), $\frac{1}{45}\sum_{j=1}^{15}\sum_{h=1}^{3}|\hat{\gamma}_{j}(h)-\gamma_{j}(h)|$, for the autocovariances. When the missing ratio is mild, for instance $\rho\leq 0.3$, iSSA, Kalman smoothing, and linear interpolation perform reasonable well, while TWI and $k$-TWI are quite close to the best-performing methods. Mean imputation significantly distorts the estimates. However, when $\rho>0.3$, TWI and $k$-TWI show some advantages. This is expected because when the missing ratio is high, accurate estimation of the missing data is difficult, but distribution matching is still possible. Table 4 shows the MAE for estimating the autocorrelations. For $\rho\leq 0.2$, most methods yield similar estimates. When $\rho>0.2$, $k$-TWI outperforms the competing methods, among which the Kalman smoothing is a strong contender. Meanwhile, TWI is somewhat close to Kalman smoothing for $\rho\leq 0.4$. Linearly interpolated series, however, has autocorrelations quite different from the original series. Finally, Table 5 shows the MAE for the partial autocorrelations. For $\rho\geq 0.3$, TWI and $k$-TWI consistently show good performance. The MAE of Kalman smoothing is similar to the proposed methods only when $\rho\leq 0.4$. Unlike in previous cases, the partial autocorrelations of the iSSA imputations are inaccurate, showing that it often neglects stochastic dynamic relationships.
Table 3: Mean absolute errors ($\times 10$) for estimating autocovariances. Mean stands for imputation by the sample mean. For each missing ratio $\rho$, the smallest two values are in boldface.
$\rho$ Linear iSSA Kalman ScalarF Mean TWI $k$-TWI
0.1 0.14 0.16 0.14 0.31 0.98 0.16 0.18
0.2 0.21 0.20 0.38 0.72 1.91 0.27 0.34
0.3 0.41 0.31 0.30 1.01 2.46 0.35 0.33
0.4 0.50 0.49 0.43 1.58 3.10 0.46 0.46
0.5 0.77 0.63 0.60 2.35 3.54 0.56 0.52
0.6 0.98 0.91 1.25 2.90 3.84 0.99 0.79
Table 4: Mean absolute errors ($\times 10$) for autocorrelations. See Table 3 for notations.
$\rho$ Linear iSSA Kalman ScalarF Mean TWI $k$-TWI
0.1 0.18 0.17 0.12 0.15 0.63 0.13 0.14
0.2 0.27 0.25 0.28 0.35 1.37 0.23 0.26
0.3 0.55 0.40 0.36 0.51 1.76 0.31 0.28
0.4 0.73 0.47 0.42 0.72 2.37 0.45 0.38
0.5 1.17 0.72 0.56 1.22 2.85 0.65 0.44
0.6 1.52 0.92 1.09 1.81 3.17 1.23 0.76
Table 5: Mean absolute errors ($\times 10$) for the partial autocorrelations. See Table 3 for notations.
$\rho$ Linear iSSA Kalman ScalarF Mean TWI $k$-TWI
0.1 1.10 1.13 1.11 1.34 2.27 1.14 1.27
0.2 1.09 1.26 1.21 1.72 2.95 1.39 1.47
0.3 1.26 1.60 1.50 1.91 3.19 1.34 1.34
0.4 1.38 2.00 1.54 2.01 3.60 1.51 1.55
0.5 1.48 1.76 1.88 2.24 3.81 1.47 1.46
0.6 1.64 2.20 2.21 2.85 4.07 1.60 1.58
Second, we use the first 200 observations from the imputed series to estimate an AR(15) model, where the estimated coefficients are then used in predicting, in a one-step-ahead rolling window fashion, the rest 108 observations from the true data. Figure 4 plots the root mean squared prediction errors. For most series except Series 1 and 2, TWI and $k$-TWI are among the methods yielding the lowest prediction errors. Their advantage is again more pronounced when the missing ratio is large. In addition, the TWI-type methods are stable as $\rho$ increases, whereas the iSSA and Kalman smoothing may suffer from large errors.
In summary, the above analysis shows that the proposed TWI and $k$-TWI methods, through distribution matching, are useful in imputing missing values of time series for downstream statistical analysis. Compared to the existing nonparametric iSSA method and the parametric Kalman smoothing, such advantage is especially valuable when the missing ratio is high.
Figure 4: Root mean squared prediction errors for each series. The AR(15) model is trained using the first 200 observations of the imputed series, which is then tested on the rest 108 data from the uncontaminated true series (excluding missing values).
7 Conclusion
This paper introduced a temporal Wasserstein imputation, a novel nonparametric method for imputing missing data in time series. Compared to existing approaches, TWI (1) does not require any model specification prior to imputation; (2) can be carried out in a principled manner, and (3) yields imputations whose distributional properties are more aligned with those of the underlying process, making downstream analyses more reliable. In simulation study and real data analysis, TWI has shown its versatility and contributions to the existing literature, especially for the cases of high missing ratio and of imputing stochastic, nonlinear time series.